OpenStudy (anonymous):

5+x/5=-2-x/2 show me the steps

OpenStudy (anonymous):

get all the x's on one side and the numbers on the other: 5 + x/5 = -2 - x/2 subtract 5 from both sides x/5 = -7 - x/2 add x/2 to both sides x/5 + x/2 = -7 simplify the fractions 2x/10 + 5x/10 = -7 7x/10 = -7 multiply both sides by 10 7x = -70 divide both sides by 7 x = -10